### v1

@@ Appendix D Sticking coefficient
[…]
@@
Since the exact value and rate of change with temperature of the sticking coefficient $\alpha$ remains highly uncertain (see the discussion in Zhukovska et al. 2016), we adopt a constant value of $\alpha=1/3$ for our unresolved high-gas density regime driven by turbulence (see Section 3.4). For comparison, the turbulence-driven subgrid model for dust accretion is driven for $\lambda_{\rm J}<4\Delta x$ and the effective temperature-density relation in our simulations is $T\simeq 10^{4}(n/0.1\,\rm H\,cm^{-3})^{-2/3}$ (see Fig. 9) for $n>0.1\,\rm H\,cm^{-3}$ and $T<10^{4}\,\rm K$, hence, the temperature at which the subgrid model is triggered for a fiducial resolution of $\Delta x=20\,\rm pc$ is $T\simeq 750\,\rm K$. This temperature gives a value of $\alpha$ from Le Bourlot et al. (2012) that is close to our adopted constant value of $\alpha$ in the subgrid model. We have run a G10LG simulation with an increased value of $\alpha=1$ in that regime: the DTM increases by $5\%$, mostly driven by an increase in the fraction of carbonaceous grains, which goes from a fiducial value of $f_{\rm C}=0.34$ to $f_{\rm C}=0.38$, and produces a significant decrease of the depletion factor of C by a factor of 2 at $n\simeq 10\,\rm H\,cm^{-3}$.
The turbulence-driven subgrid model for gas accretion onto dust grains has a maximum density above which the accretion of new refractory elements is impossible (either $n_{\rm max}=10^{3}$ or $10^{4}\,\rm H\,cm^{-3}$ (see section 3.4). This value modifies the behaviour of the effective accretion timescale as a function of the turbulent Mach number and is shown in Fig. 27 for the case of accretion onto carbonaceous dust grains. The effective time scale for accretion $t_{\rm acc,eff}$ exhibits a decrease with turbulent Mach number above $\mathcal{M}>1$ until it reaches a minimum for strong $\mathcal{M}\simeq 5-10$ after which the effective timescale increase. The exact behaviour is driven by the maximum value $n_{\rm max}$. To guide the eye, we show the corresponding free-fall time of the gas at the given mean gas densities $\bar{n}$, and the (single) time delay for SNII adopted in this work.
@@
Figure 27: The effective accretion rate of refractory materials onto grains (here for carbonaceous grains) as a function of the turbulent Mach number for two different mean density $\bar{n}=1$ and $100\,\rm H\,cm^{-3}$ (respectively in solid blue and in dashed blue). The various colors stand for different density cut-offs $n_{\rm max}=10^{3}$ and $10^{4}\,\rm H\,cm^{-3}$ (respectively in dark and light blue). For comparison we show the free-fall time corresponding to the quoted gas density $\bar{n}$ in black, and the SN time delay in red dot-dashed.

### v2

@@ Appendix D Sticking coefficient
    $\alpha_{\rm C}(T)=0.95(1+\beta T/T_{0})/(1+T/T_{0})^{\beta}$, $\beta=2.22$ and $T_{0}=56$ K measured from laboratory experiments for chemisorption of H2 molecules on silicate surfaces (Chaabouni et al. 2012).
@@
Since the exact value and rate of change with temperature of the sticking coefficient $\alpha$ remains highly uncertain (see the discussion in Zhukovska et al. 2016), we adopted a constant value of $\alpha=1/3$ for our unresolved high-gas density regime driven by turbulence (see Section 3.4). For comparison, the turbulence-driven subgrid model for dust accretion is activated for $\lambda_{\rm J}<4\Delta x$ and the effective temperature-density relation in our simulations is $T\simeq 10^{4}(n/0.1\,\rm H\,cm^{-3})^{-2/3}$ (see Fig. 9) for $n>0.1\,\rm H\,cm^{-3}$ and $T<10^{4}\,\rm K$, hence, the temperature at which the subgrid model is triggered for a fiducial resolution of $\Delta x=20\,\rm pc$ is $T\simeq 750\,\rm K$. This temperature gives a value of $\alpha$ from Le Bourlot et al. (2012) that is close to our adopted constant value of $\alpha$ in the subgrid model. We ran a G10LG simulation with an increased value of $\alpha=1$ in that regime: the DTM increases by $5\%$, mostly driven by an increase in the fraction of carbonaceous grains, which goes from a fiducial value of $f_{\rm C}=0.34$ to $f_{\rm C}=0.38$, and produces a significant decrease of the depletion factor of C by a factor of 2 at $n\simeq 10\,\rm H\,cm^{-3}$.
The turbulence-driven subgrid model for gas accretion onto dust grains has a maximum density above which the accretion of new refractory elements is impossible (either $n_{\rm max}=10^{3}$ or $10^{4}\,\rm H\,cm^{-3}$ (see section 3.4). This value modifies the behaviour of the effective accretion timescale as a function of the turbulent Mach number and is shown in Fig. 27 for the case of accretion onto carbonaceous dust grains. The effective timescale for accretion $t_{\rm acc,eff}$ exhibits a decrease with turbulent Mach number above $\mathcal{M}>1$ until it reaches a minimum for strong $\mathcal{M}\simeq 5-10$ after which the effective timescale increases. The exact behaviour is driven by the maximum value $n_{\rm max}$. To guide the eye, we show the corresponding free-fall time of the gas at the given mean gas densities $\bar{n}$, and the (single) time delay for SNII adopted in this work.
@@
Figure 27: Effective accretion rate of refractory materials onto grains (here for carbonaceous grains) as a function of the turbulent Mach number for two different mean density $\bar{n}=1$ and $100\,\rm H\,cm^{-3}$ (respectively in solid blue and in dashed blue). The various colors stand for different density cut-offs $n_{\rm max}=10^{3}$ and $10^{4}\,\rm H\,cm^{-3}$ (respectively in dark and light blue). For comparison we show the free-fall time corresponding to the quoted gas density $\bar{n}$ in black, and the SN time delay in red dot-dashed.
@@ Appendix E Sizes and extinction curves for low mass galaxies
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption